OpenStudy (anonymous):

I dont know how to enter 2 ln5+ 2 ln3 into my calculator!! please help!

OpenStudy (whpalmer4):

What kind of calculator do you have?

OpenStudy (anonymous):

TI30xa

OpenStudy (anonymous):

or know how to write it another way so i can type it in differently?

OpenStudy (whpalmer4):

Looks like you should be able to enter it just as it looks: 2 LN 5 + 2 LN 3 When it doubt, you could use parentheses.

OpenStudy (whpalmer4):

What do you get when you try to enter it?

OpenStudy (anonymous):

I get 8 but i have to do more with the problem and everytime enter the answer in my math homework online it says its wrong

OpenStudy (whpalmer4):

Hmm, you're taking the natural log of 5 and 3? That's not going to give you such a nice round number.

OpenStudy (anonymous):

the problem is 2^2-x=3^5x+7

OpenStudy (whpalmer4):

What if you just try to take the natural log of 5, what do you get?

OpenStudy (whpalmer4):

\[2^2 - x = 3^{5x} + 7\]?

OpenStudy (anonymous):

its 2 to the power of 2-x = 3 to the power of 5x+7

OpenStudy (whpalmer4):

\[2^{2-x} = 3^{5x+7}\] ?

OpenStudy (anonymous):

yes

OpenStudy (whpalmer4):

It's a good news/bad news kind of situation. Good news is that you don't have to evaluate that expression you asked about. Bad news is that means you haven't done the problem correctly :-)

OpenStudy (anonymous):

lol how do I do it?

OpenStudy (whpalmer4):

What if you multiply both sides by \(2^{2-x}\)?

OpenStudy (whpalmer4):

Then you can use a trick: \[e^{\log x} = x\] and \[e^0 = 1\]

OpenStudy (whpalmer4):

Sorry, I mistyped my first suggestion: \(2^{x-2}\) is what you want to multiply both sides with.

OpenStudy (anonymous):

so \[3^{5x+7}\times2^{2-x}\]

OpenStudy (whpalmer4):

That gives us \[2^{2-x}*2^{x-2} = 2^{x-2}3^{5x+7}\]Adding the exponents on the left (common base) \[1 = 2^{x-2}3^{5x+7}\]

OpenStudy (whpalmer4):

But with the trick I mentioned, we can write that as \[1 = e^{\log 2^{x-2}} e^{\log 3^{5x+7}}\]

OpenStudy (whpalmer4):

Now we can use the property that \[\log a^n = n \log a\]and write it as \[1 = e^{(x-2)\ln 2}e^{(5x+7)\ln 3}\]and then add the exponents of a common base to get \[1 = e^{(x-2)\ln 2 + (5x+7)\ln 3}\]Do you think you can take it from there?

OpenStudy (anonymous):

yes thank you!

OpenStudy (whpalmer4):

I need to go run some errands, but post your answer and I'll have a look when I return.

OpenStudy (anonymous):

ok